# From Prompts to Protection: Large Language Model-Enabled In-Context Learning for Smart Public Safety UAV

### v1

@@ VIII Conclusion
This paper advocates integrating LLM-enabled ICL into public safety UAV to address key operational functions, such as path planning and velocity control. By leveraging ICL, UAVs can dynamically interpret task objectives, adjust behaviors, and make context-aware decisions based on natural language instructions and a few example prompts without the need for retraining. By classifying critical UAV functions and formulating corresponding task descriptions, we demonstrate how LLMs can effectively support path planning, velocity control, data collection scheduling, and power control. The proposed LLM-enabled ICL framework offers a lightweight, adaptable, and scalable approach, setting the stage for new generation of intelligent, edge-enabled public safety UAVs. We envision the application of ICL to public safety UAV swarms, along with its real-world validation and potential use in other domains.

### v2

@@ VIII Conclusion
This paper advocates integrating LLM-assisted ICL into public safety UAV to address key operational functions, such as path planning and velocity control. By leveraging ICL, UAVs can dynamically interpret task objectives, adjust behaviors, and make context-aware decisions based on natural language instructions and a few example prompts without the need for retraining. By classifying critical UAV functions and formulating corresponding task descriptions, we demonstrate how LLMs can effectively support path planning, velocity control, data collection scheduling, and power control. The proposed LLM-assisted ICL framework offers a lightweight, adaptable, and scalable approach, setting the stage for new generation of intelligent, edge-enabled public safety UAVs. We envision the application of ICL to public safety UAV swarms, along with its real-world validation and potential use in other domains.